OpenStudy (anonymous):

Find the equation of the tangent line to x^2+2y^2=3

OpenStudy (anonymous):

tangent where?

OpenStudy (anonymous):

To the equation above

OpenStudy (anonymous):

it has to be tangent at some point, if you want to get the equation of the tangent line

OpenStudy (anonymous):

I have dy/dx as being -2x/74

OpenStudy (anonymous):

oh yeah at points (1,1) missed that part

OpenStudy (anonymous):

-2x/4y (fat fingers)

OpenStudy (anonymous):

\[ x^2+2y^2=3 \] \[2x+4yy'=0\] \[y'=-\frac{x}{2y}\]

OpenStudy (anonymous):

Right, i am that far. So I should be able to plug m (x,y) into y' to get my slope, then use point slope form to get my tang line right?

OpenStudy (anonymous):

replace \(x\) and \(y\) by 1 to get the slope, which is \(-\frac{1}{2}\) now use the point slope formula with \[m=-\frac{1}{2}, x_1=y_1=1\]

OpenStudy (anonymous):

yeah, so the second part of the question is when I am checking myself, the tang line should be on the edge of the elpise right? Because the line i made passes through the elipse and not on the edge